40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 12 17:50:21 2011, josephb, Update, CDS, Worked on c1lsc, MC2 screens 
    Reply  Wed Jan 12 22:33:24 2011, kiwamu, Update, CDS, MC2 dewhitening are healthy except for UR 
Message ID: 4146     Entry time: Wed Jan 12 22:33:24 2011     In reply to: 4144
Author: kiwamu 
Type: Update 
Category: CDS 
Subject: MC2 dewhitening are healthy except for UR 

I briefly checked the MC2 analog dewhitening filters.

It turned out that the switching of the dewhitening filters from epics worked correctly except for the UR path.

I couldn't get a healthy transfer function for the UR path probably because the UR monitor at the front panel on either the AI filter or the dewhitening filter maybe broken.

Need a check again.

Quote:  #4144

Tomorrow I'll take a look at the binary outputs and see why the analog filters aren't actually changing.

ELOG V3.1.3-